# Recovery of Sparse Graph Signals

### v1

@@ VI-B2 Case 2- Erdős Réyni Graph with $70$ Nodes
    Scenario 4 (Localized Support): In this scenario, a single node is randomly selected from the Erdős Réyni graph in Fig. 7. Then, three of the node’s neighbors are randomly chosen and added to the support set. This scenario describes a situation where the non-zero elements are clustered in close proximity within the graph structure.
Figure 7: Erdős-Reńyi graph with 70 nodes. This realization was obtained using a connection probability of 0.12 between any two nodes and was used for all simulations.
Figure 8 shows the F-score as a function of the graph filter degree, where ${\Omega}_{t}$ is drawn from Scenario $3$ and Scenario $4$ with equal probability. Unlike the results in Fig. 5, none of the methods show similar results to the optimal GIC method. The GM-GIC and its correction GM-GIC¯$^{c}$ methods obtain the closest result. Closely behind is the $\overline{\text{G-BNB}}^{c}$ method, with results up to $0.1$ better than the G-BNB method. The OMP¯$^{c}$ method shows similar results to the $\overline{\text{G-BNB}}^{c}$ method for the first $3$ graph degrees, yet lower F-scores at the higher degrees. The Lasso and $\ell_{1}$-BNB methods and their corrections show competitive results only for the first $2$ graph degrees.
Figure 8: Erdős-Rényi Graph: the F-score of the different methods where ${\Omega}_{t}$ is drawn from Scenario $3$ and Scenario $4$ with equal probability.
VI-C Run-Time
In order to demonstrate the computational complexity of the different methods empirically, their associated run-time (average computation time) is evaluated using Matlab on two Intel(R) Xeon(R) CPU E5-2660 v4 @ 2.00 GHz processors. The signal support set, ${\Omega}_{t}$, in the following simulation is drawn with equal probability from Scenario $1$ and Scenario $2$. In Fig. 9(a), we present the run-time of the different algorithms versus the graph filter degree. The optimal GIC method requires the highest run-time (operating in close to $100$ seconds). The run-time of the GM-GIC and G-BNB methods increases as the graph filter degree increases, where the GM-GIC method run-time is substantially larger than the G-BNB method run time. For example, for $\Psi=5$, the GM-GIC method requires a couple of seconds while the G-BNB  method operates in $10$ milliseconds. The G-BNB method requires a lower run-time than the $\ell_{1}$-BNB method for the first $4$ graph filter degrees. The Lasso and OMP methods perform the fastest; however, as shown in Figs. 5, 6, and 8, these methods display lower F-scores in comparison to the proposed methods. Notably, the additional run-time required by the GFOC method to correct these methods is negligible for all the methods.
Figure 9: Square Grid Graph: The run-time of the different sparse support set recovery methods versus (a) the graph filter degree; (b) the SNR; and (c) the number of nodes in the graph, where ${\Omega}_{t}$ is drawn with equal probability from the selection rule in Scenario $1$ or Scenario $2$.
In Fig. 9(b), we present the run-time versus the SNR where the graph filter degree is ${\Psi}=4$. It can be seen that the run-time of the GM-GIC method significantly increases as the SNR increases. The run-time required for the remaining methods is constant w.r.t. the SNR. The G-BNB method requires a significantly lower run-time than the GM-GIC method and yields similar results to the $\ell_{1}$-BNB methods. Figure 9(c) presents the run-time as a function of the graph size, where the filter degree is ${\Psi}=4$ and the signal support set, ${\Omega}_{t}$, is chosen with equal probability from Scenarios $1$ and $2$. The run-time required for the optimal GIC methods grows exponentially with the graph (model) size. In contrast, the GM-GIC is only slightly affected, while the G-BNB  run-time is constant across the different graph sizes. The graph size also affects the Lasso and OMP methods, but both still operate significantly faster than the other methods.
@@ VII Conclusions
In this paper, we propose three sparse recovery techniques: the GM-GIC, Graph-BNB-GIC, and GFOC methods. These techniques solve the support set recovery problem for diffused sparse graph signals, which are node-domain sparse graph signals diffused over the graph by a graph filter. The proposed methods use the GIC as the optimization cost function and leverage the underlying graphical structure to reduce the computational complexity and enhance the recovery performance. In particular, the GM-GIC and Graph-BNB-GIC directly tackle the support set recovery problem by considering local graph properties and efficiently partitioning the dictionary elements and iteratively searching over candidate support sets, respectively. The GFOC method is based on correcting the solution obtained by any other support set recovery technique. In addition, we provide a theoretical analysis of the considered setting with: 1) an examination of the dictionary matrix atoms (i.e. graph filter columns) derived from the underlying graphical structure; and 2) an evaluation of the computational complexity of the proposed methods, based on the graph filter degree, the maximum degree of the graph, and sparsity restrictions. Simulations conducted over the square grid graph and the Erdős-Rényi graph reveal that our proposed GM-GIC and Graph-BNB-GIC methods achieve higher support set recovery accuracy compared to state-of-the-art methods such as OMP, Lasso, and $\ell_{1}$-BNB, while maintaining significantly lower computational overhead than the optimal GIC method. In general, the GM-GIC method outperforms the Graph-BNB-GIC method in terms of the accuracy of the support set recovery, but requires more computing time. It is also demonstrated that the GFOC method can significantly improve the support set recovery accuracy of existing methods, including the OMP, Lasso, and $\ell_{1}$-BNB methods, without imposing significant additional computation overhead. By applying the GFOC method on the Graph-BNB-GIC method, we obtained support set estimation results comparable to the GIC benchmark, with a run-time comparable to the OMP and Lasso methods. Consequently, the best performance is obtained by combining the Graph-BNB-GIC and GFOC methods.

### v2

@@ VI-B Test Case A - Synthetic Data Set
    Scenario 2 (Localized Support): In this scenario, a single node is randomly selected from the cluster. Then, ${\lvert{\Omega}\rvert}-1$ ($3$ when ${\lvert{\Omega}\rvert}=4$) of the node’s neighbors are randomly chosen and added to the support set. This scenario describes a situation where the non-zero elements are in close proximity.
The non-zero elements of the sparse graph signal, ${{\bf{x}}}$, are drawn from a standard normal distribution. The filtered signal, ${{\bf{H}}}{\bf{x}}$, is normalized to a specified snr, which is calculated as $\text{SNR}={\|{{\bf{H}}}{{\bf{x}}}\|}^{2}/({\lvert{\mathcal{V}}\rvert}\sigma_{n}^{2})$, where the noise covariance matrix is ${\bf{R}}=\sigma_{n}^{2}{\bf{I}}$, and unless otherwise specified, $\sigma_{n}=0.01$.
Figures 4.(a)-(c) present the F-score as a function of: (a) the graph filter degree for SNR$=17$ [dB] and ${\lvert{\mathcal{V}}\rvert}=140$; (b) snr for ${\Psi}=4$ and ${\lvert{\mathcal{V}}\rvert}=140$; and (c) and graph size for ${\Psi}=4$ and SNR=$20$ [dB]. Across all figures, the gmgic and its correction gmgic¯$^{c}$ methods obtain the closest result to the gic benchmark, showing the highest F-scores. Close behind is the $\overline{\text{G-BNB}}^{c}$ method, with results up to $0.1$ better than the G-BNB method in Fig. 4.(a). The results also show that gfoc significantly enhances the $\ell_{1}$-BnB method across all experiments.
In Fig. 4.(a), it can be seen that the F-score of all methods decreases almost always as the graph filter degree increases. The omp¯$^{c}$ method shows similar results to the $\overline{\text{G-BNB}}^{c}$ method for the first $2$ graph degrees, yet lower F-scores at the higher degrees. The Lasso and $\ell_{1}$-bnb methods and their corrections show competitive results only for the first $2$ graph degrees. In Fig. 4.(b) it can be seen that the F-score of all methods increases as the SNR increases. The corrected version of the omp method, omp¯$^{c}$, improves the omp F-score by approximately $25\%$ compared to the original omp. Similarly, the corrected version of the $\ell_{1}$-bnb, $\overline{\ell_{1}-BnB}^{c}$ method, improves the F-score of the $\ell_{1}$-bnb by up to $20\%$. However, it remains less competitive than the omp¯$^{c}$, $\overline{\text{G-BNB}^{c}}$, and gmgic methods. The Lasso method and its corrected version exhibit poor F-score performance across all observed snr.
Figure 4.(c) displays the F-score as a function of graph size (number of sbm clusters). For ${\lvert{\mathcal{V}}\rvert}=70$, the graph contains one cluster. When ${\lvert{\mathcal{V}}\rvert}=140$, it has two clusters, with the signal support drawn from Cluster $1$. For ${\lvert{\mathcal{V}}\rvert}=210$, the graph has three clusters, with the support either entirely from Cluster $1$ or uniformly from Clusters $1$ and $3$. In all cases, the support follows Scenario $2$, and the snr is set to SNR=$20$ dB. It can be seen that the graph size has a minimal impact on the F-score of the proposed methods. The main effect is seen with the $\ell_{1}$ and Lasso methods (and their corrections), which perform worse compared to the other approaches.
Figures 4.(d)-(f) presents the mse of the different methods as a function of the (a) graph filter degree, (b) the snr, and (c) the graph size, using the same settings as those used for the F-score evaluation in Figs. 4.(a)-(c). The mse performance in Figs. 4.(d)-(f) aligns closely with the F-scores results in Figs. 4.(a)-(c), leading to similar conclusions regarding the comparative performance of the different methods and the superiority of the proposed methods. For example, the gic, gmgic, and $\overline{\text{G-BNB}^{c}}$ methods, which achieve the highest F-scores in Figs. 4.(a)-(c), also exhibit the lowest mse in Figs. 4.(d)-(f). In addition, it can be seen in Fig. 4.(e) that the snr significantly impacts the performance of the gic, gmgic, and $\overline{\text{G-BNB}^{c}}$ methods, showing improvements of approximately $0.7$ across the tested range. In contrast, the omp method improves by only $0.1$ and shows no further improvement beyond an snr of $8$ [dB]. Similarly, omp¯$^{c}$ improves by around $0.4$ and shows no further improvement beyond an snr of $14$ [dB].
In Figs. 4.(g)-(i) we present the run-time of the different methods as a function of the (g) graph filter degree, (h) the snr, (i) and the graph size, respectively, for the same setting as used for the F-score and mse evaluation in Figs. 4.(a-c) and Figs. 4.(d)-(f). Across all figures, the gic benchmark method requires the highest run-time, exceeding $1,000$ seconds per Monte-Carlo simulation when the graph size is greater than $140$ nodes. In contrast, the proposed gmgic and G-BNB methods have significantly reduced run-times while still maintaining good F-scores and mse results, as shown in Figs. 4.(a)-(c) and Figs. 4.(d)-(f). The $\ell_{1}$-bnb, Lasso, and omp methods are the fastest overall. However, as shown in Figs. 4.(a)-(c) and Figs. 4.(d)-(f), these methods display lower F-scores and mse, respectively, than the proposed methods. Notably, the additional run-time required by the gfoc method to correct these methods is negligible.
In Fig. 4.(g) it can be seen that the run-time of the gmgic and G-BNB methods increases as the graph filter degree increases, where the gmgic method run-time is larger than the G-BNB method run-time. In Fig. 4.(h), it can be seen that the run-time of the gmgic method significantly increases as the snr increases. The run-time of the G-BNB method moderately increases and does not change for snr higher than $15$ [dB]. Both methods still operate significantly faster than the gic method. The run-time required for the remaining methods is constant wrt the snr and lower than the methods mentioned above. In Fig. 4.(i) it can be seen that the run-time required for the gic method grows exponentially with the graph (model) size. In comparison, the gmgic run-time grows significantly between ${\lvert\mathcal{{\mathcal{V}}}\rvert}=70$ to ${\lvert\mathcal{{\mathcal{V}}}\rvert}=140$, but remains constant between ${\lvert\mathcal{{\mathcal{V}}}\rvert}=140$ to ${\lvert\mathcal{{\mathcal{V}}}\rvert}=210$. The G-BNB  method is only marginally affected by graph size. In addition, the graph size also affects the run-time of the Lasso and omp methods, where their run-time becomes closer to those of the bnb methods as $|\mathcal{V}|$ increases. To conclude, the results demonstrate that the proposed gmgic and gbnb scale well with the graph size and enable efficient processing even in larger network configurations.
VI-C Test Case B - Graph Blind Deconvolution of Brain Signals
Graph blind deconvolution has various real-world applications [49, 48]. In particular, it has been used for the analysis of epileptic seizures [79, 80] considered in this subsection. In this context, the graph signals are the neural activity levels in regions of interest (ROIs) within the brain, and the graph nodes and edges represent the brain regions and the anatomical connection between them, respectively. This application has been explored in gsp blind deconvolution studies [49, 48].
In order to apply our graph-based sparse recovery methods to the problem of blind deconvolution, we propose the following two-stage approach. In the first stage, a blind deconvolution is performed to estimate the unknown graph filter coefficients, $\hat{h}_{i}$, $i=0,\ldots,{\Psi}$, from the input measurements, ${{\bf{y}}}$. This estimation is implemented by one of the existing blind deconvolution methods, such as the methods in [49, 48]. The graph filter can then be recovered based on (6), where the gso associated with the underlying graph is assumed to be known. Then, in the second stage, our graph-based sparse recovery methods are employed to reconstruct the sparse graph signal ${{\bf{x}}}$ based on the model in (11), where the graph filter used is the one estimated in the previous stage. This overall two-stage approach is illustrated in Fig. 5.
Blind Deconvolution Graph-Based Sparse Recovery $\hat{{{\bf{x}}}}$${{\bf{y}}}$$\{\hat{h}_{k}\}_{k=1}^{{\Psi}}$
Figure 5: Two-stage approach for blind deconvolution combined with graph-based sparse recovery: The input measurements ${{\bf{y}}}$ undergo a blind deconvolution process to estimate the graph filter coefficients, $\{\hat{h}\}_{k}$, $k=1,\ldots,{\Psi}$. Then, a graph-based sparse recovery is performed by the proposed methods, resulting in the estimated signal $\hat{{{\bf{x}}}}$.
(a) Known filter, F-score
(b) Blind Nuc. norm, F-score
(c) Blind Logdet, F-score
(d) Known filter, mse
(e) Blind Nuc. norm, mse
(f) Blind Logdet, mse
Figure 7: Brain networks: F-score and mse of graph-based sparse recovery methods using graph coefficients that are either known (see (a) and (d)) or estimated via the blind deconvolution techniques: Blind Nuc. norm ((b) and (e)) and Blind Logdet ((c) and (f)).
In the following simulations, we consider the model in (11), where the graph filter is unknown and the underlying graph is a brain network. We use the Laplacian gso, which is computed based on an adjacency matrix drawn from one of the six brain networks provided in [62, 81]. Entries in each row of the Laplacian that are significantly smaller (i.e. two orders of magnitude smaller than the maximum entry in the same row) are set to zero, and the Laplacian is subsequently normalized such that its largest eigenvalue is $12$ to establish a controlled setting for the analysis.
The graph filter is constructed by setting the graph filter degree to ${\Psi}=3$, substituting the Laplacian matrix and the graph filter coefficients (which are uniformly drawn from the interval $[0,1]$) into (6), and normalizing the result to achieve a Frobenius norm of $1$. The signal support is set to ${\Omega}=4$ and is drawn according to Scenario $2$ from Subsection VI-B. The non-zero values of the graph signals are uniformly drawn from the interval $[0,1]$ and normalized to a unit-norm vector. This setup follows a similar framework to that considered in [49].
For the first stage of graph filter estimation, we use the following blind deconvolution techniques:
  • Blind Nuc. norm: Employs nuclear norm minimization combined with $\ell_{1}$ row sparsity regularization (see Equation (9) in [49]).
  • Blind Logdet: Employs a log surrogate combined with $\ell_{2}$ row sparsity regularization (see Equation (17) with $\tau_{x}=0.5$ and $\tau_{h}=0$ in [48]).
The performance of each scenario is evaluated through at least $300$ Monte Carlo simulations.
We first evaluate the existence of Assumption A.1, which assumes a low maximal degree for the graph, in the context of the real-data brain networks. This assumption inherently indicates that the graph is sparse. Table I presents the ratio of the number of edges to the number of possible connections, $\frac{{\lvert{\mathcal{E}}\rvert}}{\frac{1}{2}{\lvert{\mathcal{V}}\rvert}({\lvert{\mathcal{V}}\rvert}-1)}$, and the maximal degree, $d_{\text{max}}$, for the six real-data brain networks derived from the dataset in [67]. as described above. This table demonstrates that these brain networks are sparse graphs with low maximal degrees, thus aligning with Assumption A.1.
$\mathcal{G}_{1}$ $\mathcal{G}_{2}$ $\mathcal{G}_{3}$ $\mathcal{G}_{4}$ $\mathcal{G}_{5}$ $\mathcal{G}_{6}$
$\frac{{\lvert\text{edges}\rvert}}{\frac{1}{2}{\lvert{\mathcal{V}}\rvert}({\lvert{\mathcal{V}}\rvert}-1)}$ 0.1 0.11 0.09 0.11 0.11 0.11
$d_{\text{max}}$ 3.28 4.41 5.83 3.87 3.93 3.68
TABLE I: Brain networks: the ratio of the number of edges and the number of possible connections, $\frac{{\lvert\text{edges}\rvert}}{{\lvert K_{n}\rvert}}$, and the maximal degree, $d_{\text{max}}$, for the real-data brain networks from [67, 81].
Next, we evaluate the graph-based methods from Subsection VI-A for three scenarios: 1) using a known graph filter; and 2) and 3) using the output of the two-stage blind deconvolution approach with the two different techniques, referred to as “Blind Nuc. norm” and “Blind Logdet”, respectively. We compare the result of the different methods with those of the existing blind deconvolution methods, which can be interpreted as performing only Stage 1 of the approach described in Fig. 5. These methods are referred to as “Blind” in Fig. 7. In particular, in Figs. 5.(b) and 5.(e), the Blind Nuc. norm method is used, while in Figs. 5.(c) and 5.(f) the BLind Logdet method is employed. For the clarity of presentation, we only present the results for the graph-based methods after applying the gfoc method. This is because the gfoc method consistently improves the recovery performance with negligible additional computational overhead (run-time).
Figure 7 presents the F-score and mse performance of the different methods as a function of the snr. The mse is computed as $MSE=\frac{1}{{\|{{\bf{x}}}{\bf{h}}^{T}\|}}{\|{{\bf{x}}}{\bf{h}}^{T}-\hat{{{\bf{x}}}}\hat{{\bf{h}}}^{T}\|}_{F},$ where ${{\bf{x}}}$ is the sparse graph signal, $\hat{{{\bf{x}}}}$ is its estimator, ${\bf{h}}$ represents the vector of graph filter coefficients, and $\hat{{\bf{h}}}$ is its estimator. For a known graph filter, we use $\hat{{\bf{h}}}={\bf{h}}$.
Figures 7.(b)-(c) and 7.(e)-(f) show that the gic, gmgic¯$^{c}$, $\overline{\text{G-BNB}}^{c}$, and omp¯$^{c}$ methods significantly outperform the blind deconvolution techniques in both F-score and mse under the case where the graph filter is unknown. These results confirm that the proposed sparse recovery methods are also applicable in scenarios where the graph filter is not known a priori. In addition, it can be seen in these figures that while omp¯$^{c}$ achieves competitive results at low snr levels, its performance degrades as the snr increases, and is worse than those of the gmgic¯$^{c}$ and $\overline{\text{G-BNB}}^{c}$ methods. The $\overline{\text{Lasso}}^{c}$, is competitive to the above methods only at high snr values, and the $\overline{\ell_{1}\text{-}BnB}^{c}$ method works poorly for all snr rates and even downgrades the blind deconvolution output. Notably, the F-scores of the graph-based sparse recovery methods, using the two-stage approach from Fig. 5, are comparable to those obtained where the graph filter is known, as can be seen by comparing the results in Figs. 7(a) with the results in Figs. 7. (b)-(c). In contrast, the mse is more sensitive to the accuracy of the graph filter coefficients, indicating that improving the filter estimation could lead to better mse results. It is also important to emphasize that the additional run-time for the $\overline{\text{G-BNB}}^{c}$, omp¯$^{c}$, $\overline{\ell_{1}\text{-}BnB}^{c}$, and $\overline{\text{Lasso}}^{c}$ methods have been found to be negligible compared to the computational cost of the first stage of the blind deconvolution. Thus, it is recommended that a subsequent graph sparse recovery stage be employed on the existing graph blind deconvolution methods.
VI-D Test Case C - Cholera Outbreak
In this section, we evaluate the performance of the different methods in a real-world test case where both the graph (network) and the graph filter are unknown. While the observation model in (11) forms the basis for our approach and underpins the proposed sparse recovery methods, it is important to note that in this scenario, the true structure of the real-world data and the underlying model are unknown. Thus, the simulations in this subsection demonstrate how the proposed framework can be applied to the practical application of interpreting real-world data, despite the uncertainty in the true graph structure and graph filter that govern the system dynamics.
The real-world data for this case study, drawn from [68], involves the 1854 Cholera outbreak in London’s Soho district, where Dr. John Snow identified the source of the epidemic as a contaminated water pump [69]. Inspired by the framework in [40], this study applies the proposed graph-based sparse recovery methods to identify the source of the outbreak using historical data. The used dataset includes information on death tolls across 250 buildings, their geographic locations, and the positions of 8 water pumps, as provided in [68].
First, we model the underlying graph using a type-aware nearest neighbor structure. The graph is constructed based on the hypothesis that water was the primary transmission vector for cholera, with nodes representing buildings and pumps. For the graph edges, there is an edge between each pump and its two nearest pumps, each pump and its five nearest buildings, and each building and its two nearest buildings. The edge weights are computed using a Gaussian kernel:
$W_{k,m}=\exp\left(-\frac{(\text{dist}(k,m))^{2}}{2\cdot(f(k,m))^{2}}\right),$
where dist represents the Euclidean distance, and $f(i,j)$ depends on the node types: $f(k,m)=1000$ for pump connections, $f(k,m)=80$ for pump-building connections, and $f(k,m)=20$ for buildings connections. For ground truth, we consider the infected pump and the building with the highest death toll. To promote sparsity, we reduce the death count by one in each building, resulting in $136$ non-zero elements out of $258$, and we normalize the data vector to $1$. It should be noted that the underlying graph and the graph filter are set to satisfy Assumptions A.1-A.2. Using the reconstructed graph, we apply the methodology outlined in Fig. 5, which includes a blind deconvolution step followed by graph-based sparse recovery. This follows the approach detailed in Subsection VI-C, with the estimated graph. Accordingly, we use the same methods as those described in Subsection VI-C, but with the estimated graph structure. The performance is evaluated using at least $60$ Monte Carlo simulations.
Figure 8: Cholera network: the F-score of graph-based sparse recovery methods using coefficients estimated via the Bilnd Nuc. norm method.
Figure 8 shows the F-scores of the graph-based sparse recovery methods: gic, gmgic¯$^{c}$, $\overline{\text{G-BNB}}^{c}$, omp¯$^{c}$, $\overline{\ell_{1}\text{-}BnB}^{c}$, and $\overline{\text{Lasso}}^{c}$, as a function of snr. In each simulation, we added a zero-mean Gaussian additive noise to the data vector to obtain the specific snr. This evaluation focuses on the challenge of identifying the ground truth provided by the data vector contaminated with noise under different snr. The results indicate that the blind deconvolution method Blind Nuc. norm [49] cannot reconstruct the support set at any snr level in this case. In contrast, the gic, gmgic¯$^{c}$, $\overline{\text{G-BNB}}^{c}$, and omp¯$^{c}$ methods have a significantly higher F-score and achieve close to perfect reconstruction (F-score$>0.9$) for large snr. These methods achieve similar F-scores for most snr, where the F-score increases as the snr increases. Between $15$ [dB] and $25$ [dB], $\overline{\text{G-BNB}}^{c}$ achieves slightly higher results. Thus, the results in Fig. 8 demonstrate that the proposed graph-based sparse recovery methods offer substantial improvements in support set recovery, even in challenging real-world scenarios with unknown graph structures and filters, outperforming the existing blind deconvolution techniques. Notably, the run-time added by the graph-based methods, excluding the gic method, is negligible in comparison to the run-time of the blind deconvolution method.
@@ VII Conclusions
In this paper, we propose three sparse recovery techniques: the gmgic, gbnb, and gfoc methods. These techniques solve the support set recovery problem for diffused sparse graph signals, which are node-domain sparse graph signals diffused over the graph by a graph filter. The proposed methods use the gic as the optimization cost function and leverage the underlying graphical structure to reduce the computational complexity and enhance the recovery performance. In particular, the gmgic and gbnb directly tackle the support set recovery problem by considering local graph properties and efficiently partitioning the dictionary elements and iteratively searching over candidate support sets, respectively. The gfoc method is based on correcting the solution obtained by any other support set recovery technique. Furthermore, we provide a theoretical analysis of the considered setting with: 1) an examination of the dictionary matrix atoms (i.e. graph filter columns) derived from the underlying graphical structure; and 2) an evaluation of the computational complexity of the proposed methods, based on the graph filter degree, the maximum degree of the graph, and sparsity restrictions.
Simulations conducted over the sbm graphs reveal that the proposed gmgic and gbnb methods achieve higher support set recovery and signal recovery accuracy compared to state-of-the-art methods such as omp, Lasso, and $\ell_{1}$-bnb, while maintaining significantly lower computational overhead than the optimal gic method. In general, the gmgic method outperforms the gbnb method in terms of the accuracy of the support set recovery, but requires more computing time. It is also demonstrated that the gfoc method can significantly improve the support set recovery accuracy of existing methods, including the omp, Lasso, and $\ell_{1}$-bnb methods, without imposing significant additional computation overhead. By applying the gfoc method on the gbnb method, we obtained support set estimation results comparable to the gic benchmark, with a run-time comparable to the omp and Lasso methods. Consequently, the best performance is obtained by combining the gbnb and gfoc methods. In addition, we show the application of these methods for the problem of blind deconvolution in which the graph signal and the graph filter are unknown. Simulations conducted on brain networks and real data on the Cholera $1854$ outbreak show that applying the proposed methods significantly improves the signal recovery performance for the application of blind deconvolution.